### v2

# A Multi-Task and Multi-Label Classification Model for Implicit Discourse Relation Recognition

Abstract
We address the inherent ambiguity in Implicit Discourse Relation Recognition (IDRR) by introducing a novel multi-task classification model capable of learning both multi-label and single-label representations of discourse relations. Our model is trained exclusively on the DiscoGeM corpus and evaluated both on the DiscoGeM and the PDTB 3.0 corpus. We establish the first benchmark on multi-label IDRR classification and achieve SOTA results on single-label IDRR classification using the DiscoGeM corpus. Finally, we present the first evaluation on the potential of transfer learning between the DiscoGeM and the PDTB 3.0 corpus on single-label IDRR classification.
A Multi-Task and Multi-Label Classification Model
for Implicit Discourse Relation Recognition
Nelson Filipe Costa  and Leila Kosseim Computational Linguistics at Concordia (CLaC) Laboratory Department of Computer Science and Software Engineering Concordia University, Montréal, Québec, Canada nelsonfilipe.costa@mail.concordia.ca leila.kosseim@concordia.ca
1 Introduction
Discourse analysis is responsible for understanding the relations of coherence between segments of a text. While different frameworks have been proposed for the analysis of discourse, one of the most widely used is the Penn Discourse Treebank (PDTB) (Miltsakaki et al., 2004; Prasad et al., 2008). In the PDTB framework, a discourse relation is established between two textual segments (called arguments) either explicitly, through the use of a discourse connective, or implicitly when there is no connective. The sense of the relation is then determined from a list of 36 senses organized hierarchically in three levels (Prasad et al., 2019). The first level has four broad senses and the following levels have subsequently more detailed senses. At the implicit level, the recognition of the sense of the relation - known as implicit discourse relation recognition (IDRR) - remains arguably the hardest task in discourse analysis. This comes as a result of the inherently subjective nature of the task, where even trained human annotators find it difficult to agree on the sense annotation of implicit discourse relations (Rohde et al., 2016; Scholman and Demberg, 2017; Hoek et al., 2021).
In fact, such ambiguity is present in most tasks where language is concerned, since different language constructs can carry different meanings or interpretations (Poesio, 2020; Basile et al., 2021; Jiang and de Marneffe, 2022). Yet, most NLP classification tasks are established on the idea that a single gold-label should exist and neglect that there could be genuine human variation in labeling due to disagreement, subjective interpretation or the existence of multiple plausible answers (Pavlick and Kwiatkowski, 2019; Uma et al., 2021; Plank, 2022). In the field of discourse, the idea that discourse annotation can often be ambiguous is not new (Stede, 2008) and has been recently highlighted by the difficulties in establishing a clear mapping between implicit discourse relations across frameworks (Demberg et al., 2019; Costa et al., 2023).
For this reason, a growing body of work has recently considered disagreement between humans annotators not as noise, but as an opportunity to enhance the accuracy and relevance of NLP models by capturing the complexity of human interpretations and biases (Nie et al., 2020; Uma et al., 2020; Fornaciari et al., 2021; van der Meer et al., 2024). This new perspective has lead to a recent paradigm shift in implicit discourse annotation. For instance, rather than relying on expert annotators to find a single gold label for each implicit relation, the DiscoGeM corpus (Scholman et al., 2022a) crowdsourced this task to multiple laypeople to obtain a distribution of labels for each relation following the PDTB 3.0 framework (Prasad et al., 2019).
This paper addresses the inherent ambiguity in IDRR by proposing novel multi-task implicit discourse relation classification model capable of learning both multi-label and single-label representations of discourse relations. Our model is trained exclusively on the DiscoGeM corpus (Scholman et al., 2022a) and evaluated both on the DiscoGeM and the PDTB 3.0 corpus (Prasad et al., 2019). Overall, the main contributions of this work are:
  • We present the first fully multi-label approach to IDRR by training a multi-task classification model on the DiscoGeM corpus to produce probability distributions across multiple labels for all instances. Our model can then be further adapted to the traditional single-label task by selecting the label with the highest probability from the generated distribution.
  • We establish the first benchmark on multi-label IDRR classification and achieve SOTA results on single-label IDRR classification using the DiscoGeM corpus.
  • We present an in-depth analysis on the potential of transfer learning between the DiscoGeM and the PDTB 3.0 corpus on single-label IDRR classification.
2 Previous Work
The advent of transformer-based models (Vaswani et al., 2017) has lead to a significant improvement in performance across all NLP tasks and in IDRR in particular. While traditional model would rely on recurrent neural networks (Shi and Demberg, 2019a; Liang et al., 2020), recent models leverage the power of pre-trained large language models (LLMs) through fine-tuning (Long and Webber, 2022; Liu and Strube, 2023; Wu et al., 2023) or prompt-tuning (Zhou et al., 2022; Zhao et al., 2023; Chan et al., 2023) to achieve SOTA results. However, the SOTA performance of these models has only slightly increased over the last years.
One possible explanation for this lack of significant improvement can be attributed to the fact that current approaches to IDRR rely on the principle that there is a single ground truth in the interpretation of implicit discourse relations (Aroyo and Welty, 2013, 2015; Dumitrache et al., 2019). However, IDRR is an inherently ambiguous task, often with possible multiple interpretations (Rohde et al., 2016; Scholman and Demberg, 2017; Hoek et al., 2021). Thus, a new line of work has recently tried to capture this source of ambiguity by pioneering new discourse annotated corpora capturing a distribution of sense labels per relation instead of a single reference label (Yung et al., 2019; Pyatkin et al., 2020; Scholman et al., 2022a, b; Pyatkin et al., 2023). Following this line of work, recent works have leveraged the potential of such multi-labelled discourse relations for IDRR (Yung et al., 2022; Pyatkin et al., 2023; Costa and Kosseim, 2024). However, even if trained on multi-labelled data, these models still predict a single-label sense per relation. On the other hand, Long et al. (2024) leveraged the $4.9\%$ of implicit discourse relations in the PDTB 3.0 (Prasad et al., 2019) annotated with two senses to develop classification models capable of predicting up to two senses per relation. Yet, to the best of our knowledge, no previous work has truly captured the full multi-label distribution of all possible senses across all instances.
3 Our Approach
In this work we propose a novel multi-task classification model capable of simultaneously predicting probability distributions across all three sense levels in the PDTB 3.0 framework. Figure 1 shows the architecture of our model. The model takes as input a concatenated pair of discourse arguments (ARG1+ARG2) and generates an embedding using the pre-trained RoBERTa (Liu et al., 2019) language model [^1] We also experimented with other pre-trained models, but RoBERTa achieved the best performance (see Section 5.1).. The embedding is then fed to a linear and dropout layer, before being passed to three separate classification heads - one for each of the three sense levels. In the multi-label setting, the output of each classification head is a probability distribution over the existing senses at the corresponding level. In the single-label setting, an additional pooling layer selects the sense with the highest probability in the corresponding distribution as the output of each classification head. We use the Adam optimization algorithm (Kingma and Ba, 2015) to minimize the loss function of our model, which we calculate as the sum of the individual losses of each classification head.
Figure 1: The model takes as input a pair of discourse arguments and generates one output for each sense level in the PDTB 3.0 framework. In the multi-label setting each output is a vector of probabilities over the existing senses at each level, while in the single-label setting each output is the sense (selected by the pooling layers in faded colors) with the highest probability score.
Multi-Label Classification.
To calculate the loss of each multi-label classification head, we use the mean absolute error (MAE) loss function (see Equation 2 in Appendix B) as it leads better performance (see Section 5.1). Similarly to other multi-label classification works in NLP (Pyatkin et al., 2023; van der Meer et al., 2024), we evaluate our model in the multi-label classification setting using the Jensen-Shannon (JS) distance (Lin, 1991) to measure the similarity between the predicted and the target probability distributions.
Single-Label Classification.
To calculate the loss of each single-label head, we used the cross-entropy (CE) loss function (see Equation 1 in Appendix B) as it emphasizes the correct classification of the highest probability sense label (see Section 5.1). To evaluate our model in the single-label classification setting, we use the weighted F1-score between the labels with the highest scores in the predicted and the target distributions.
4 Data Preparation
We used two different discourse annotated corpora: DiscoGeM (Scholman et al., 2022a) and the PDTB 3.0 (Prasad et al., 2019). Our model was trained exclusively on the DiscoGeM corpus and evaluated on both DiscoGeM (for multi- and single-label classification) and on the PDTB 3.0 (for single-label classification). Although they use different annotation methodologies, both corpora follow the same discourse framework (Webber et al., 2019). This enabled our model to leverage the knowledge acquired from DiscoGeM and directly apply it to the PDTB 3.0 through transfer learning.
4.1 DiscoGeM
The DiscoGeM corpus (Scholman et al., 2022a) contains 6,807 inter-sentential implicit discourse relations distributed across four textual genres: 2,800 relations from political texts, 3,060 from literary texts, 645 from encyclopedic texts and 302 relations extracted from the PDTB 3.0 corpus (Prasad et al., 2019). Each relation was annotated by at least 10 different laypeople. These annotations were then averaged to produce a multi-label sense distribution across 29 different senses in the PDTB 3.0 framework - the Belief and the Speech-Act level-2 senses were not annotated in DiscoGeM.
To use DiscoGeM, we adapted the standard set of 14 level-2 senses proposed by Kim et al. (2020) for IDRR classification in the PDTB 3.0. Since DiscoGeM did not include the Cause+Belief sense from the standard set, and to maintain a 14 label set, we replaced it by the Similarity sense as it was the most frequent of the remaining level-2 senses [^2] We could not replace Cause+Belief by another Contingency sense, because they were all already considered.. In addition to this set of level-2 senses, we also considered level-3 senses when possible. When no level-3 sense was available, we selected the corresponding level-2 sense as replacement. Table 8 in Appendix A shows the distribution of senses across all levels in the original DiscoGeM and in our adapted set of senses.
Figure 2: Distribution of majority label senses across all splits of the DiscoGeM corpus.
For single-label classification, we replace the full multi-label sense distribution of each discourse relation in DiscoGeM by its sense with the highest score - referred to as the majority label. To reduce computational complexity and maintain consistency across model evaluations, we used the same training, validation and testing splits in both the multi and single-label settings. We split 70% of the corpus for training, 10% for validation and 20% for testing. To ensure a balanced distribution, we kept the same distribution of majority labels across all data splits as shown in Figure 2. We opted for a fixed training, validation and test split, instead of multiple fold splits for cross-validation, to streamline the experimentation process as we conducted an extensive number of experiments.
4.2 PDTB 3.0
The majority of research in IDRR relies on the single-label annotated PDTB 3.0 corpus (Prasad et al., 2019). The PDTB 3.0 corpus was annotated by expert annotators and consists of 53,631 discourse relations extracted from Wall Street Journal news articles - from which 21,827 are implicit. Since there are currently no other benchmarks in multi-label IDRR, we also evaluated our model on the traditional single-label classification task using different test splits of the PDTB 3.0 corpus.
To allow transfer learning between DiscoGeM and the PDTB 3.0, we used the same set of 14 level-2 senses described in Section 4.1. Following the common approach to IDRR, with the PDTB 3.0 we only kept the level-1 and level-2 senses. To compare our work against SOTA models in single-label IDRR, we generated the two commonly used Lin (Lin et al., 2009) and Ji (Ji and Eisenstein, 2015) test splits - the former uses section 23 in the PDTB 3.0 corpus as the test set, while the latter uses sections 21-22. However, since recent works highlighted the limitations of using such small test sets to draw meaningful conclusions (Shi and Demberg, 2017; Kim et al., 2020), we also generated test splits following the cross-validation scheme proposed by Kim et al. (2020). Table 9 in Appendix A shows the total number of level-2 sense instances on the different single-label test sets.
4.3 Corpora Agreement
Since 302 discourse relations from the PDTB 3.0 were co-annotated in DiscoGeM, we calculated the annotation agreement between both corpora. Table 1 shows the number of relations where the reference label from the PDTB 3.0 agreed with the majority label (top-1) of DiscoGeM, or was found among the top-3, top-5 or top-10 labels in DiscoGeM. The fact that for 87 (18.8%) of the jointly annotated relations, the reference label from the PDTB 3.0 corpus was not selected by at least one of the annotators of DiscoGeM highlights the inherent ambiguity in IDRR.
Top-1 Top-3 Top-5 Top-10
$92\;(30.5\%)$ $164\;(54.3\%)$ $197\;(65.2\%)$ $215\;(71.2\%)$
Table 1: Number of instances where the reference label in the PDTB 3.0 was found within the top-k labels in DiscoGeM for the set of 302 co-annotated relations.
5 Experiments
In order to optimize performance in terms of lower JS distances for the multi-label classification and in terms of higher weighted F1-scores for the single-label classification, we evaluated different configurations of our model (see Fig. 1) on the DiscoGeM validation set (see Fig. 2). All of the reported results are the average scores of three different runs with different random starts for 10 epochs and with a batch size of 16. We ran all of our experiments on a 32-core compute node with 512GB of RAM and all of the code used can be found on GitHub [^3] https://anonymous-link.
5.1 Model Selection
LM Loss JS Distance (Multi-Label) F1-Score (Single-Label)
Level-1 Level-2 Level-3 Level-1 Level-2 Level-3
BERT MAE $0.314\pm 0.004$ $0.461\pm 0.007$ $0.535\pm 0.004$ $63.34\pm 1.41$ $46.44\pm 0.82$ $39.20\pm 1.82$
CE $0.328\pm 0.004$ $0.563\pm 0.002$ $0.634\pm 0.004$ $63.49\pm 1.38$ $51.42\pm 0.58$ $47.65\pm 0.92$
DistilBERT MAE $0.330\pm 0.006$ $0.472\pm 0.003$ $0.543\pm 0.003$ $61.29\pm 0.96$ $39.29\pm 0.66$ $31.75\pm 0.55$
CE $0.341\pm 0.005$ $0.571\pm 0.006$ $0.639\pm 0.003$ $60.28\pm 0.58$ $45.55\pm 1.52$ $41.19\pm 1.50$
RoBERTa MAE $0.304\pm 0.010$ $0.448\pm 0.006$ $0.530\pm 0.004$ $65.12\pm 1.20$ $51.81\pm 1.54$ $42.99\pm 1.09$
CE $0.317\pm 0.003$ $0.561\pm 0.003$ $0.630\pm 0.005$ $65.48\pm 0.82$ $53.51\pm 0.39$ $49.55\pm 0.11$
DistilRoBERTa MAE $0.318\pm 0.001$ $0.463\pm 0.003$ $0.539\pm 0.004$ $61.30\pm 0.38$ $42.75\pm 1.63$ $34.96\pm 1.29$
CE $0.329\pm 0.002$ $0.564\pm 0.005$ $0.636\pm 0.003$ $65.02\pm 0.69$ $52.83\pm 0.71$ $47.94\pm 0.88$
Table 2: Results of experimenting with different pre-trained language models and different loss functions in multi-label classification (JS distance) and in single-label classification (weighted F1-score). The results were averaged across three different runs with random starts. Values in bold show the best score for each metric.
Starting with the claim that LMs pre-trained with a next sentence prediction (NSP) objective lead to better results in IDRR (Shi and Demberg, 2019b), we evaluate the performance of using BERT (Devlin et al., 2019) as the backbone of our model against RoBERTa (Liu et al., 2019), which was used in recent works on single-label IDRR (Long and Webber, 2022; Liu and Strube, 2023) and was not pre-trained on NSP. We then experimented with distilled versions of both models (Sanh et al., 2019). Finally, we evaluated the performance of each LM with respect to the MAE and the CE loss functions as shown in Table 2.
Table 2 shows that for both loss functions the RoBERTa model outperforms BERT both in terms of lower JS distances in the multi-label setting and higher F1-scores in the single-label setting across all sense levels. This shows that the benefit of using larger models (RoBERTa against BERT) outweighs previous findings that LMs pre-trained with the NSP training objective lead to better results in IDRR (Shi and Demberg, 2019b). Following this observation, both RoBERTa and BERT models outperform their distilled versions across all metrics. Finally, Table 2 also shows that the choice of loss function plays an important role in multi- and single-label IDRR. Training the model with the MAE loss leads to lower JS distances in all configurations, while training the model with the CE loss leads to better F1-scores. Therefore, we chose RoBERTa as the backbone of the model and use the MAE loss when training for multi-label IDRR and the CE loss when training for single-label IDRR. We also experimented with the mean squared error (MSE) and the Huber loss. However, they performed slightly worse as shown in Table 10 in Appendix B.
5.2 Fine-Tuning
LR Decay JS Distance (Multi-Label with MAE) F1-Score (Single-Label with CE)
Level-1 Level-2 Level-3 Level-1 Level-2 Level-3
$1e^{-4}$- $0.386\pm 0.004$ $0.509\pm 0.003$ $0.569\pm 0.005$ $36.07\pm 1.67$ $21.61\pm 0.58$ $10.05\pm 0.99$
$5e^{-5}$- $0.339\pm 0.041$ $0.478\pm 0.039$ $0.554\pm 0.016$ $64.71\pm 2.39$ $49.76\pm 1.65$ $45.52\pm 1.59$
[…]
Linear $0.305\pm 0.009$ $0.459\pm 0.007$ $0.547\pm 0.004$ $64.70\pm 0.35$ $54.32\pm 2.15$ $48.96\pm 0.81$
CosAn $0.299\pm 0.004$ $0.447\pm 0.008$ $0.529\pm 0.008$ $65.15\pm 0.61$ $53.99\pm 1.17$ $49.56\pm 1.25$
[…]
Linear $0.315\pm 0.006$ $0.480\pm 0.007$ $0.566\pm 0.007$ $64.84\pm 1.14$ $55.86\pm 1.10$ $50.34\pm 1.55$
CosAn $0.306\pm 0.003$ $0.469\pm 0.002$ $0.558\pm 0.004$ $65.40\pm 1.04$ $54.97\pm 1.85$ $49.62\pm 1.26$
[…]
Table 3: Results of experimenting with different learning rates and decay functions in multi-label classification (JS distance) using the MAE loss and in single-label classification (weighted F1-score) using the CE loss. The results were averaged across three different runs with random starts. Values in bold show the best score for each metric.
We fine-tuned our model separately with different learning rates for multi-label IDRR and for single-label IDRR. Table 3 shows the results of fine-tuning our model with the MAE loss in the multi-label setting and with the CE loss in the single-label setting. Under both settings, there is a significant increase in performance when training with a learning rate of $lr=1e^{-5}$ and $lr=5e^{-6}$. For this reason we further experimented training with the linear and cosine annealing decay (Loshchilov and Hutter, 2017) functions for these two learning rates. The linear decay function was set to reduce the learning rate during the first 5 epochs to a final value of half of its original value and the cosine annealing decay function was set to vary between the original learning rate and half of its value for two complete periods during the 10 epochs. Based on the results in Table 3, the warm restart mechanism of the cosine annealing decay function with a learning rate $lr=1e^{-5}$ achieved the best results in the multi-label setting and the linear decay function with a learning rate $lr=5e^{-6}$ achieves the best results in the single-label setting.
6 Results
Table 4 shows the results of our multi- and single-label models on the test set of DiscoGeM, while Table 5 shows the transfer learning results of our single-label model on the different test splits (see Section 4.2) of the PDTB 3.0. The only difference between the multi- and single-label models is that the first was trained with the MAE loss and the second was trained with the CE loss (see Section 3).
6.1 Results on DiscoGeM
Models JS Distance (Multi-Label) F1-Score (Single-Label)
Level-1 Level-2 Level-3 Level-1 Level-2 Level-3
Baseline $0.519\pm 0.002$ $0.636\pm 0.001$ $0.714\pm 0.002$ $34.72\pm 1.63$ $22.42\pm 1.34$ $18.81\pm 0.80$
[…]
 Costa and Kosseim (2024) -- -- -- -- -- $51.38\pm n/a$
[…]
Ours w/ CE $0.323\pm 0.003$ $0.564\pm 0.003$ $0.634\pm 0.002$ $65.89\pm 1.35$ $55.99\pm 1.73$ $50.82\pm 1.26$
Table 4: Final results on the test set of DiscoGeM in multi-label classification (JS distance) and in single-label classification (weighted F1-score). The results were averaged across three different runs with random starts. Values in bold show the best score for each metric.
Since there is no previous work on full multi-label IDRR, we created a random predictor to use as baseline. The baseline predicts a probability distribution over all the senses at each sense level by emulating the annotation protocol of DiscoGeM. The baseline makes ten single-label predictions following the probability distribution of each sense at each level and the final probability distribution is computed as the average of these ten predictions. In the single-label setting, we compared our results against previous work on single-label IDRR using DiscoGeM (Yung et al., 2022; Costa and Kosseim, 2024).  Yung et al. (2022) trained and tested their model on both DiscoGeM and the PDTB 3.0 on single-label classification only at level-2. When using DiscoGeM they adapt it to the single-label setting by generating ten single-label instances for each discourse relation proportional to its probability distribution in the corpus.  Costa and Kosseim (2024) follow a similar approach to ours in the sense that they also train their model on multi-label discourse relations in DiscoGeM to generate a probability distribution from which they select the majority sense as their predicted single-label.
Table 4 shows that our model is consistently better than the baseline on all metrics and significantly outperforms the current SOTA in single-label level-2 classification (Yung et al., 2022) ($55.99$ vs $23.66$). At level-3, our model achieves results at par with SOTA (Costa and Kosseim, 2024) taking the standard deviation into account ($50.82\pm 1.26$ vs $51.38$). Furthermore, it is important to highlight that the model of Costa and Kosseim (2024) was specifically trained to predict senses at level-3, while our proposed model can generalize across all levels.
6.2 Transfer Learning Results on PDTB 3.0
Models F1-Score (PDTB 3.0)
Lin Ji Cross
Level-1 Level-2 Level-1 Level-2 Level-1 Level-2
 Yung et al. (2022) -- -- -- $38.07\pm 2.25$- --
 Long and Webber (2022) -- -- $70.05\pm n/a$ $57.62\pm n/a$- --
 Liu and Strube (2023) -- -- $71.15\pm 0.47$ $54.92\pm 0.81$- --
[…]
Ours w/ CE $50.44\pm 0.81$ $34.38\pm 0.86$ $49.43\pm 1.47$ $33.11\pm 1.57$ $49.72\pm 1.45$ $33.21\pm 1.67$
Table 5: Transfer learning results on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). The Lin and Ji results were averaged across three different runs with random starts, while the Cross results were averaged across all 12 folds. Values in bold show the best score for each metric.
The results of transfer learning from DiscoGeM to the PDTB 3.0 on the single-label classification are shown in Table 5. Our results are compared against the work of Yung et al. (2022) and the SOTA (Long and Webber, 2022; Liu and Strube, 2023). Both SOTA approaches used a RoBERTa based model and were evaluated using the Ji split (Ji and Eisenstein, 2015) and the cross-validation scheme proposed by Kim et al. (2020).
Table 5 shows that our model has a lower performance than previous work on all metrics for which there was a reference result [^4] We could not find any work on single-label IDDR using the Lin split for the PDTB 3.0 corpus.. However, it is important to note that our model was never exposed to any PDTB 3.0 data in its training. The lower performance in this setting can further be explained by the agreement across corpora shown in Table 1 in Section 4.3, where the majority sense in DiscoGeM matched only on $92(30.5\%)$ of the co-annotated discourse relations in the PDTB 3.0. This shows that a direct transfer learning approach is not possible between both corpora. This also explains the contrasting performances shown in Tables 4 and 5 between our model and the one from Yung et al. (2022), which was trained on the two corpora. Our model performs significantly better when evaluated on DiscoGeM ($55.99$ vs $23.66$), but worst when evaluated on the PDTB 3.0 ($33.11$ vs $38.07$). In the future, after training on DiscoGeM, we could subsequently fine-tune our model on the PDTB 3.0 to improve its performance in this setting.
Level-1 DiscoGeM PDTB 3.0 Level-2 DiscoGeM PDTB 3.0
Test Lin Ji Cross Test Lin Ji Cross
Temporal $61.27\pm 1.34$ $4.32\pm 1.58$ $15.75\pm 4.50$ $16.23\pm 4.86$ Synchronous $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Asynchronous $57.87\pm 0.84$ $6.01\pm 2.05$ $22.02\pm 4.42$ $23.16\pm 6.65$
Contingency $58.67\pm 1.83$ $36.12\pm 7.17$ $34.27\pm 8.24$ $41.08\pm 2.87$ Cause $65.36\pm 0.12$ $47.21\pm 1.93$ $46.40\pm 2.47$ $48.93\pm 2.07$
Condition -- $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Purpose -- $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Comparison $39.31\pm 3.15$ $32.65\pm 1.69$ $37.67\pm 5.20$ $33.72\pm 2.34$ Concession $36.24\pm 4.55$ $29.70\pm 5.09$ $27.92\pm 1.75$ $19.46\pm 3.70$
Contrast $4.13\pm 2.94$ $11.10\pm 7.92$ $12.70\pm 9.35$ $10.43\pm 6.61$
Similarity $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Expansion $75.39\pm 1.15$ $64.44\pm 0.66$ $62.69\pm 0.38$ $63.48\pm 1.16$ Conjunction $52.56\pm 0.66$ $42.30\pm 1.26$ $39.91\pm 1.97$ $44.45\pm 3.04$
Equivalence -- $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Instantiation $44.84\pm 0.18$ $12.15\pm 9.12$ $17.90\pm 12.72$ $15.62\pm 10.49$
Level-of-Detail $46.85\pm 3.28$ $33.61\pm 2.90$ $29.32\pm 2.70$ $29.27\pm 3.06$
Manner -- $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Substitution $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Table 6: Individual per-class results for each sense in level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification (weighted F1-score). Senses marked with -- were not present in the test set. The DiscoGeM Test and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
7 Analysis
To further evaluate the performance and robustness of our model, we present two additional analyses. First, we conduct a per-class evaluation to assess the ability of the model in classifying each individual sense at level-1 and level-2. Second, we investigate the coherence of the predictions by evaluating how well the predicted labels at level-1 and level-2 align with the hierarchical structure of senses of the PDTB 3.0 framework. For easier interpretation, we consider only single-label results.
7.1 Per-Sense Results
To identify disparities in performance between individual labels and to assess how well the model is able to generalize across senses, we evaluate the single-label predictions of our model at level-1 and level-2. Table 6 shows the weighted F1-score of each predicted sense in the test set of DiscoGeM and in the different test sets of the PDTB 3.0.
The per-sense performance of the model on the test set of DiscoGeM at level-1 and level-2 in Table 6 follows the distribution of senses in the corpus (see Table 8 in Appendix A). Note that the (Condition, Purpose, Equivalence and Manner) senses never occur as the majority label of a relation in DiscoGeM (see Figure 2 in Section 4.1) and are therefore not present in the test set (see Table 9 in Appendix A). Table 6 shows that our model has difficulties predicting under-represented senses in the corpus, such as Synchronous, Contrast, Similarity and Substitution. One possible solution would be to use data augmentation techniques for these under-represented instances. The confusion matrices in Figures 3 and 4 in Appendix C provide further details into the per-sense performance on the test set of DiscoGeM. The performance on the test sets of the PDTB 3.0 shown in Table 6 is lower with respect to the performance on DiscoGeM, similar to what was shown in Table 4 in Section 6.1 and in Table 5 in Section 6.2. This drop in performance is more pronounced for some senses, such as Temporal at level-1 and Asynchronous and Instantiation at level-2. Conversely, the model performs better at the level-2 Contrast sense. Since the model was not trained on any PDTB 3.0 data, these discrepancies can be explained by inconsistent annotations across both corpora as highlighted in Table 1 in Section 4.3.
7.2 Sense Coherence across Levels
Level-1 DiscoGeM PDTB 3.0 Level-2 DiscoGeM PDTB 3.0
Test (%) Lin (%) Ji (%) Cross (%) Test (%) Lin (%) Ji (%) Cross (%)
Temporal $92.94\pm 2.93$ $93.27\pm 6.31$ $90.93\pm 4.51$ $92.48\pm 4.93$ Synchronous $n/a$ $n/a$ $n/a$ $n/a$
Asynchronous $90.00\pm 1.44$ $82.74\pm 6.76$ $82.44\pm 11.34$ $88.30\pm 8.08$
Contingency $98.91\pm 0.20$ $100.00\pm 0.00$ $99.74\pm 0.45$ $100.00\pm 0.00$ Cause $62.37\pm 4.92$ $38.19\pm 6.45$ $38.42\pm 8.27$ $29.93\pm 1.32$
Condition -- $n/a$ $n/a$ $n/a$
Purpose -- $n/a$ $n/a$ $n/a$
Comparison $86.31\pm 1.22$ $73.81\pm 13.21$ $68.09\pm 10.47$ $58.45\pm 4.12$ Concession $81.19\pm 2.83$ $89.15\pm 8.10$ $89.36\pm 8.98$ $97.93\pm 1.52$
Contrast $91.67\pm 14.43$ $90.77\pm 10.09$ $90.77\pm 11.11$ $92.68\pm 8.19$
Similarity $n/a$ $n/a$ $n/a$ $n/a$
Expansion $70.70\pm 1.26$ $37.91\pm 10.82$ $65.02\pm 12.30$ $53.01\pm 2.21$ Conjunction $97.61\pm 0.71$ $97.14\pm 0.03$ $3.20\pm 0.88$ $96.69\pm 0.96$
Equivalence -- $n/a$ $n/a$ $n/a$
Instantiation $100.00\pm 0.00$ $100.00\pm 0.00$ $100.00\pm 0.00$ $100.00\pm 0.00$
Level-of-Detail $99.82\pm 0.31$ $99.80\pm 0.35$ $100.00\pm 0.00$ $100.00\pm 0.00$
Manner -- $n/a$ $n/a$ $n/a$
Substitution $n/a$ $n/a$ $n/a$ $n/a$
Table 7: Percentage of the instances that were classified with coherent senses at level-1 and level-2 on the test set of DiscoGeM and on different test splits of the PDTB 3.0 in single-label classification. Senses marked with -- were not present in the test set and senses marked with $n/a$ were never predicted by the model. The DiscoGeM Test and the PDTB 3.0 Lin and Ji results were averaged across three different runs with random starts, while the PDTB 3.0 Cross results were averaged across all 12 folds.
To assess the coherence of the predictions of the model across sense levels, Table 7 shows the percentage of times a sense at level-1 was predicted with a coherent sense at level-2 (and vice-versa) in the test set of DiscoGeM and in the different test sets of the PDTB 3.0. This enables us to examine the extent to which the predictions of the model are consistent with the hierarchical sense structure defined by the PDTB 3.0 framework.
As Table 7 shows, the Contingency sense at level-1 is very often predicted with a coherent level-2 sense (i.e., Cause, Condition or Purpose). Similarly, the level-2 senses Instantiation, Level-of-Detail and Conjunction are often predicted with a coherent level-1 sense. However, other senses, such as Comparison and Expansion at level-1 and Cause at level-2, are more often predicted with a contradicting sense. A possible explanation might be that these senses are inherently hard to distinguish and the model is biased towards the most represented sense. For instance, the level-2 senses Cause and Asynchronous are one of the most often co-annotated pair of senses in the DiscoGeM and Cause appears approximately three times more often than Asynchronous (see Table 8 in Appendix A), the model might be more biased towards Cause when distinguishing between these two senses. One possible method to generate more coherent predictions across sense levels, would be to input the predicted sense distributions of classification head 1 into the other two classification heads (see Figure 1), as well as input the predicted sense distributions of classification head 2 into classification head 3.
8 Conclusion
In this work we presented a novel approach to one of the most challenging tasks in discourse analysis: IDRR. We trained a multi-task classification model on DiscoGeM to learn both multi-label and single-label representations of discourse relations. Our model establishes a first benchmark on multi-label IDRR and achieves SOTA results in single-label IDRR using DiscoGeM (see Section 6.1). Our single-label classification model was also evaluated on different test splits of the PDTB 3.0 using transfer learning where it achieved lower performance than existing work (see Section 6.2). These results show that direct transfer learning between both corpora still remains a challenge.
Future research should explore the application of data augmentation techniques, such as paraphrasing, to determine whether augmenting under-represented senses in the DiscoGeM could enhance the performance of the model on individual senses (see Section 7.1). Additionally, future research should investigate whether cascading information from higher-level to lower-level classification heads within the model could improve the coherence of its predictions across sense levels (see Section 7.2). Lastly, it would be worthwhile to examine whether a model trained to learn multi-label representations of discourse relations in DiscoGeM could be be further fine-tuned on the PDTB 3.0 to achieve superior performance in single-label classification using its corpus.
9 Limitations
There are a few limitations to our work. The first limitation comes from the fact that the Cause+Belief sense was not annotated in DiscoGeM. This forced us to adapt the standard set of 14 level-2 senses proposed by Kim et al. (2020) which is used in other works on IDRR using the PDTB 3.0. To overcome this issue and maintain a 14-label set, we proposed to replace the Cause+Belief level-2 sense with the Similarity level-2 sense. However, both senses belong to different level-1 senses. Ideally, we would have replaced the Cause+Belief sense with another sense under the Contingency level-1 sense, but they were all already considered. So, we replaced it with the Similarity sense which was the most frequent sense of the remaining level-2 senses.
The second limitation of our work comes from the choice of pre-trained language model to generate the embeddings of the pair of discourse arguments. Due to computational and time limitations, we could not explore the use of larger language models, such as one of the recently released Gemini 1.5 models (Reid et al., 2024). This could have substantially increased the performance of our models. Nevertheless, we still acknowledge that the extensive experimental analysis conducted in this work would not have been possible without access to a high-performance computing facility. Finally, the use of these powerful resources also entails not negligible carbon footprint which we did not monitor during any of our experiments in this work.
Appendix A Data Statistics
To standardize and ensure a fair comparison of results on IDRR, Kim et al. (2020) proposed a set of 14 level-2 senses in the PDTB 3.0 framework. However, the Cause+Belief sense in this set was not annotated in DiscoGeM. Therefore, we adapted the standard set to accommodate the existing level-2 senses in the corpus by replacing Cause+Belief with Similarity. We then removed all senses not included in the standard set from the corpus and normalized the distribution values over the remaining senses following the L1 norm. This ensures the sum of each distribution adds to 1 for each instance while preserving the relative distance between senses. Table 8 shows the distribution of senses across all levels in the original DiscoGeM corpus and in our adapted set of 14 level-2 senses. Each value represents the sum of the corresponding sense across all instances. We then split 70% of the corpus for training, 10% for validation and 20% for testing. To ensure a balanced distribution, we kept the same distribution of majority labels across all data splits as shown in Figure 2 in Section 4.1. In the single-label setting, we replaced the full multi-label sense distribution of each discourse relation by its majority sense. Table 9 shows the level-2 majority senses in the test set of DiscoGeM and reference senses in the different test splits of the PDTB 3.0 (see Section 4.2).
Level-1 Sum Level-2 Sum Level-3 Sum
Original Adapted Original Adapted Original Adapted
Temporal $584.3$ $619.5$ Synchronous $95.8$ $102.7$
Asynchronous $488.5$ $516.8$ Precedence $448.1$ $474.2$
Succession $40.5$ $42.7$
Contingency $1,745.1$ $1,822.9$ Cause $1,740.2$ $1,819.0$ Reason $382.5$ $400.1$
Result $1,357.7$ $1,418.9$
NegResult $0.0$ $0.0$
Condition $1.2$ $1.2$ ARG1-as-Cond $0.1$ $0.1$
ARG2-as-Cond $1.1$ $1.1$
Neg-Condition $1.1$ $0.0$ ARG1-as-NegCond $1.0$ $0.0$
ARG2-as-NegCond $0.1$ $0.0$
Purpose $2.6$ $2.7$ ARG1-as-Goal $1.7$ $1.7$
ARG2-as-Goal $0.9$ $0.9$
Comparison $831.5$ $878.0$ Concession $517.1$ $548.7$ ARG1-as-Denier $169.3$ $179.9$
ARG2-as-Denier $347.7$ $368.8$
Contrast $202.4$ $213.4$
Similarity $112.0$ $116.0$
Expansion $3,034.6$ $3,184.6$ Conjunction $1,441.1$ $1,518.0$
Disjunction $2.9$ $0.0$
Equivalence $19.0$ $19.8$
Exception $1.4$ $0.0$ ARG1-as-Exception $0.3$ $0.0$
ARG2-as-Exception $1.3$ $0.0$
Instantiation $388.8$ $403.8$ ARG1-as-Instance $16.9$ $17.8$
ARG2-as-Instance $371.9$ $386.0$
Level-of-Detail $1,137.0$ $1,196.1$ ARG1-as-Detail $160.8$ $170.2$
ARG2-as-Detail $976.2$ $1,025.9$
Manner $4.6$ $4.8$ ARG1-as-Manner $1.3$ $1.3$
ARG2-as-Manner $3.3$ $3.4$
Substitution $39.7$ $42.1$ ARG1-as-Substitution $0.0$ $0.0$
ARG2-as-Substitution $39.7$ $42.1$
Table 8: Distribution of senses across all levels in the original DiscoGeM and in our adapted set of 14 level-2 senses (see Section 4.1). Each value represents the sum of the corresponding sense across all instances.
Level-1 Level-2 DiscoGeM PDTB 3.0
Test Lin Ji Cross
Temporal Synchronous $13$ $19$ $43$ $42.85\pm 11.94$
Asynchronous $103$ $33$ $105$ $99.54\pm 24.95$
Contingency Cause $515$ $271$ $406$ $465.38\pm 77.28$
Condition 00 $10$ $15$ $16.31\pm 3.61$
Purpose 00 $46$ $89$ $110.08\pm 22.60$
Comparison Concession $104$ $83$ $98$ $121.23\pm 22.87$
Contrast $27$ $52$ $54$ $69.92\pm 18.99$
Similarity $10$ $1$ $2$ $2.23\pm 1.48$
Expansion Conjunction $336$ $161$ $236$ $349.38\pm 86.25$
Equivalence 00 $24$ $30$ $27.54\pm 10.27$
Instantiation $53$ $70$ $124$ $118.62\pm 24.97$
Level-of-Detail $202$ $194$ $208$ $263.69\pm 26.94$
Manner 00 $4$ $17$ $17.92\pm 6.49$
Substitution $4$ $27$ $26$ $30.23\pm 8.28$
Total $1,367$ $995$ $1,453$ $1,734.92\pm 271.89$
Table 9: Distribution of level-2 majority label senses across the test set of DiscoGeM (see Section 4.1) and the different test splits of the PDTB 3.0 (see Section 4.2). The values in PDTB 3.0 Cross correspond to the averaged distribution of majority label senses across all 12 test folds.
Appendix B Loss Functions
[…]
  • Cross-Entropy (CE)
    $loss^{h}(\hat{y}^{h},y^{h})=-\frac{1}{N}\sum^{N}_{n=1}\sum^{C^{h}}_{c=1}log\Biggl{(}\frac{e^{\hat{y}^{h}_{n,c}}}{\sum^{C^{h}}_{i=1}e^{\hat{y}^{h}_{n,i}}}\Biggr{)}y^{h}_{n,c}$
  • Mean Absolute Error (MAE)
    $loss^{h}(\hat{y}^{h},y^{h})=\frac{1}{NC^{h}}\sum^{N}_{n=1}\sum^{C^{h}}_{i=1}|\hat{y}^{h}_{n,i}-y^{h}_{n,i}|$
  • Mean Squared Error (MSE)
    $loss^{h}(\hat{y}^{h},y^{h})=\frac{1}{NC^{h}}\sum^{N}_{n=1}\sum^{C^{h}}_{i=1}(\hat{y}^{h}_{n,i}-y^{h}_{n,i})^{2}$
  • Huber Loss (Huber)
    $loss^{k}(\hat{y}^{h},y^{h})=\sum^{N}_{n=1}\sum^{C^{h}}_{i=1}\begin{cases}\frac{{\delta^{h}_{n,i}}^{2}}{2NC^{h}}&\text{, if }|\delta^{h}_{n,i}|<1\\[11.38092pt] \frac{2|\delta^{h}_{n,i}|-1}{2NC^{h}}&\text{, otherwise}\end{cases}$
    with
    $\delta^{h}_{n,i}=\hat{y}^{h}_{n,i}-y^{h}_{n,i}$ (5)
The MAE and MSE losses in Equations 2 and 3, respectively, aim to minimize the overall difference between the predicted and target distributions by penalizing errors across all possible labels. MAE minimizes the average absolute differences, leading to predictions that are closer to the target distribution in an averaged sense, while MSE places a larger penalty on larger errors due to its quadratic nature, which can result in a stronger emphasis on outliers. The Huber loss in Equation 4 combines the properties of MAE and MSE in function of the delta parameter defined in Equation 5. In other words, the Huber loss behaves like MAE for smaller errors and like MSE for larger errors. The results of experimenting with the MAE, MSE and Huber losses were therefore relatively similar. However, the MAE was slightly better and thus better suited for multi-label IDRR as shown in Table 2 in Section 5.1. Table 10 shows the results of experimenting with the MSE and Huber loss functions in multi-label IDRR (JS distance) and in single-label IDRR (weighted F1-score). Conversely, the CE loss in Equation 1 focuses on maximizing the probability of the label with the highest score in the target distribution which makes it better suited for single-label IDRR as shown in Table 2 in Section 5.1.
Model Loss JS Distance (Multi-Label) F1-Score (Single-Label)
Level-1 Level-2 Level-3 Level-1 Level-2 Level-3
BERT MSE $0.319\pm 0.005$ $0.462\pm 0.002$ $0.544\pm 0.005$ $63.82\pm 0.83$ $48.68\pm 0.16$ $41.36\pm 1.55$
Huber $0.318\pm 0.007$ $0.463\pm 0.001$ $0.542\pm 0.002$ $64.36\pm 0.48$ $49.45\pm 0.96$ $41.58\pm 1.08$
DistilBERT MSE $0.336\pm 0.007$ $0.481\pm 0.003$ $0.558\pm 0.003$ $60.10\pm 0.96$ $43.77\pm 0.86$ $37.66\pm 1.14$
Huber $0.337\pm 0.002$ $0.479\pm 0.003$ $0.557\pm 0.006$ $60.41\pm 0.86$ $44.06\pm 0.64$ $37.29\pm 0.28$
RoBERTa MSE $0.308\pm 0.012$ $0.455\pm 0.001$ $0.534\pm 0.005$ $65.28\pm 0.36$ $53.94\pm 1.80$ $44.91\pm 1.31$
Huber $0.305\pm 0.007$ $0.450\pm 0.004$ $0.539\pm 0.003$ $65.13\pm 0.60$ $53.39\pm 0.19$ $44.33\pm 0.64$
DistilRoBERTa MSE $0.313\pm 0.004$ $0.470\pm 0.005$ $0.555\pm 0.003$ $64.32\pm 1.99$ $50.48\pm 0.29$ $42.43\pm 0.22$
Huber $0.318\pm 0.004$ $0.472\pm 0.001$ $0.561\pm 0.004$ $64.11\pm 0.72$ $50.46\pm 0.73$ $42.41\pm 0.30$
Table 10: Results of experimenting with different pre-trained language models and different loss functions in multi-label classification (JS distance) and in single-label classification (weighted F1-score). The results were averaged across three different runs with random starts. Values in bold show the best score for each metric.
Appendix C Confusion Matrices
To provide further details into the per-sense performance on the test set of DiscoGeM in Section 7.1, we generated a confusion matrix at level-1 and level-2 for the results of a single run on the test set of DiscoGeM as shown in Figures 3 and 4, respectively. The confusion matrix in Figure 3 shows that, with the exception with Comparison, the most predicted label per true label always aligns with the correct label. Nevertheless, the confusion matrix also shows that the model is more biased towards the most represented senses level-1 senses Expansion and Contingency (see Table 8 in Appendix A). At the level-2, the confusion matrix in Table 4 shows that the model is not able to predict less represented senses in DiscoGeM (see Table 8 in Appendix A), such as Synchronous, Similarity and Substitution. However, for the other senses, the most predicted label per true label often aligns with the correct label.
Figure 3: Confusion matrix for the individual per-class results of each sense in level-1 on the test set of DiscoGeM.
Figure 4: Confusion matrix for the individual per-class results of each sense in level-2 on the test set of DiscoGeM.